TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-08 23:36:00
subject: 2\20 Pt-1 ISS On Orbit Status 20-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\20 ISS On-Orbit Status 20-02-2003
Part 1 of 2

ISS On-Orbit Status 2/20/03

All ISS systems continue to function nominally, except as noted 
previously or below. Day 89 in space for Expedition 6.

Early in the morning, all crewmembers performed part 2 of the MedOps 
PHS (periodic health status) assessment, collecting blood samples for 
analysis with the U.S. PCBA (portable clinical blood analyzer) and the 
Russian MO-10 "Hemokrit" equipment. [Taking turns as test subject and 
CMO (crew medical officer), the crew was free to do these activities 
in whatever order they chose, as long as they were completed before 
breakfast and before the first exercise session. While PCBA analyzes 
total blood composition, MO-10 particularly measures the hematocrit 
(red blood cell mass) value of the blood (as a well-known phenomenon 
of space flight, red blood cell mass {normal range: 30-45%} tends to 
go down over time). Later in the day, the crew also took the PHS 
subject evaluation exam. All data, including from MO-10, PCBA, PHS, 
and yesterday's MO-9 urinalysis, were entered in the IFEP (in-flight 
examination program) on the MEC (medical equipment computer).  After 
the PHS data registration, FE-2/SO Don Pettit stowed all hardware.]

R3 Transition: After a one-day delay, uploading of the new CCS 
(command & control software) revision R3 began at about 9:00am EST, 
involving over 1100 CCS and DDCT (display definition configuration 
table) files uplinked last week to SSMMUs (solid-state mass memory 
units). [Loading was executed in a specific sequence of steps: first, 
CCS R3 into the standby and backup C&C MDMs (command & control 
multiplexer/demultiplexers); second, INTSYS R2 into the backup INT 
MDM; third, GNC (guidance, navigation & control) R3 into the backup 
GNC MDM; and fourth, version 12A.011 s/w into two PCS (personal 
computer system) laptops, which were then to be rebooted by the crew. 
The actual step-up (transition) to the new software is scheduled for 
tomorrow, with Saturday (2/22) as a backup day. All issues from 
Tuesday's aborted upload, generally associated with checksum errors 
caused by new file sizes, were verified via ground testing, are fully 
understood and were corrected.  Ground testing has also identified the 
problem with the failed PCS software load as a simple hang-up due to 
an undisplayed prompt, which was easily corrected.]

FE-1 Nikolai Budarin continued preventive maintenance on the Service 
Module (SM) air ventilation system, today cleaning Group A fans and 
grilles.

Budarin then deactivated the gas analyzer (GA) in the FGB module and
exchanged its carbon dioxide (CO2) filter assembly (BF) with a new 
unit. GA was reactivated and the spent BF stowed for disposal.

Afterwards, Nikolai moved to the Soyuz TMA-1 crew return vehicle, 
checked out the ventilation system air duct in the Descent Module to 
verify airflow, then activated the Soyuz gas analyzer. It will be left 
on for the next four days and then powered off on 2/24 (Monday).

Since Vozdukh and MCA (major constituents analyzer) are working 
nominally and the flow between modules appears to be more uniform 
since the recent IFM (in-flight maintenance), Flight Surgeon and ECLSS 
(environment control and life support system) personnel have deleted 
the requirement for the twice-daily CO2 sampling by the crew with the 
CDMK (CO2 monitoring kit).

After Don Pettit was unable on 1/17, after multiple attempts, to 
complete the TOCA (total organic carbon analyzer) water sampler 
reagent mixing, engineers and toxicology experts have now determined 
that the TOCA is unrepairable and no longer viable. A new TOCA will be 
flown as manifest space allows, but for the remainder of Expedition 6, 
the crew will not be scheduled for any further TOCA activities.

CDR Ken Bowersox and Don Pettit conducted a scheduled inventory audit 
in the Joint Airlock (A/L). [Specific tasks included locating and 
inventorying the contents of CTBs (cargo transfer bags), inspecting 
various EVA equipment and tools, reporting on available number of 
fluid QD SPDs (quick disconnect spool positioning devices), inspecting 
PRDs (payload retention devices), inventorying EVA lights, etc.]

In the Lab module, Bowersox collected samples from unknown 
debris/deposits discovered on the starboard CCAA (common cabin air 
assembly) air conditioner vent and on the GLA (general luminaire 
assembly) lighting fixture, for subsequent analysis on the ground. The 
deposits were then cleaned from all surfaces with wipes and deionized 
water.

Pettit prepared the daily IMS database update file for 
export/downlink, and Bowersox performed the regular routine 
maintenance of the SOSh life support system, while Budarin completed 
the periodic inspection of the BRPK air/condensate separator of the 
SRV-K water processing system.

Budarin conducted a consumables audit of the Russian SGO health 
maintenance system, verifying availability and location of a large 
number of EDV water container and KTO solid waste container 
components, as well as individual trash bags.

At 9:50am EST, the crew participated in two live TV media interviews, 
with National Public Radio (NPR, Scott Simon) and CBS Radio Network 
(Peter King, Bill Harwood), about nine minutes each per client. The 
station residents again were outstanding.

As an item on their task list, the crew was asked, at their leisure, 
to provide an inventory of two TVIS treadmill kits, both holding 
assorted spare parts.

A conjunction (close encounter) with the TRMM satellite (Tropical 
Rainfall Measuring Mission, object #25063) is predicted for tomorrow 
night (Friday, 2/21). Time of closest approach (TCA) is 9:25 pm EST. 
With a currently estimated overall miss distance of 14.12 km, there is 
no threat at this time and no avoidance maneuver required. Will be 
monitored closely through TCA.

Today's targets for the CEO (crew earth observations program) were 
Eastern Mediterranean Dust and Smog (a strong storm center was moving 
into the western Med. By the time of this pass, storm winds should 
have picked up dust over Algeria, Tunisia, and Libya. As the crew 
approached the Libyan coast, they were to try for oblique views left 
of track), Eastern Mediterranean Dust and Smog (this second pass over 
northern Italy and the Adriatic Sea offered a particularly good 
opportunity to document smog over the industrialized Po River Valley, 
looking left of track as ISS approached the Italian coast), Chicago, 
Illinois (crew was to use this pass over "Chicagoland" to look left of 
track up Lake Michigan and its eastern shore for ice buildup in bays 
and lake-effect snow patterns), Salt Lake City, Utah (Salt Lake City 
was at nadir this pass, but the crew may have gotten some good 
mountain snow pack views of the Wasatch Range either side of track and 
the compact Uinta Mountains to the right of track), and Tessina, Italy 
(although Tessina was probably not visible, this pass offered yet 
another look at the smog of the Po River Valley, but this time to the 
right of track).

CEO images can be viewed at the website http://eol.jsc.nasa.gov

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.